OpenStudy (ranga):

Integration of trig function with a large even exponent. Problem in comment:

OpenStudy (ranga):

\[\int \frac{(cosec^2(x)-2014)}{cos^{2014}(x)}dx\]

OpenStudy (freckles):

sec?

OpenStudy (freckles):

oh csc?

OpenStudy (freckles):

I'm not used to the above abbreviation for cosecant sorry

OpenStudy (ranga):

Even with answer choices this one is not easy to figure out. \[A) \Large -sec^{2015}(x)cot(x) + sin(x) + C\\ B) \Large -sec^{2014}(x)cot(x) + C\\ C) \Large -cosec^{2014}(x)tan(x) + C\\ D) \Large -sec^{2014}(x)cosec^2(x) + C\]

OpenStudy (freckles):

Well you could find the derivative of your answer choices to see which one gives you the expression in the integral.

OpenStudy (yttrium):

do you really need it now? i'm too sleepy now will check it tomorrow.

OpenStudy (freckles):

let me see if i can actually integrate this sucker

OpenStudy (ranga):

First the derivatives have to be taken and then they have to be rearranged to see if they can be expressed in the form of the original expression. What puzzles me is that the constant 2014 does not appear as a factor in any of the answers.

OpenStudy (freckles):

i think the sub tan(x/2)=u might really help

OpenStudy (freckles):

if i didn't make an error, it seems to be working out very pretty like

OpenStudy (ranga):

And you are getting one of the answers?

OpenStudy (freckles):

still simplifying

OpenStudy (freckles):

I got this \[\frac{u^5}{10}+\frac{u^3}{3}-\frac{2014u^3}{3}+\frac{1}{2}u-2014u+C \text{ where } \tan(x/2)=u\] so far seems ugly

OpenStudy (freckles):

I wonder if I made a mistake because I don't have any really really large exponents

OpenStudy (ranga):

Something ain't right. There should be at least one large exponent in the neighborhood of 2014.

OpenStudy (freckles):

oops I see what i did i mistaked (1-u^2) with (1+u^2) looked at the signs too fast when simplifying

OpenStudy (freckles):

so we have this with tan(x/2)=u so far: \[2 \int\limits_{}^{}\frac{(u^2+1)^3(1+u^2)^{2014}-4u^2 \cdot 2014 (u^2+1)^{2015}}{4u^2(1-u^2)^{2014}} du\]

OpenStudy (freckles):

which i don't see how we could make this pretty

OpenStudy (ranga):

The first two expressions in the numerator can be simplified to (u^2+1)^2017

OpenStudy (freckles):

well i know that but i'm saying after the easy stuff lol

OpenStudy (ranga):

Out of the trigonometric frying pan into the algebraic fire.

OpenStudy (freckles):

I will be right back. I have to eat. I will think about the problem.

OpenStudy (ranga):

Okay, thanks.

OpenStudy (ranga):

\[ \int \frac{cosec^2(x)-2014}{cos^{2014}(x)}dx = \int \frac{cosec^2(x)}{cos^{2014}(x)}dx - 2014\int \frac{1}{cos^{2014}(x)}dx = \\ \int \frac{1}{sin^2(x) * cos^{2014}(x)}dx - 2014\int \frac{1}{cos^{2014}(x)}dx = \\ u = tan(\frac x2); ~~~~dx = \frac{2du}{1+u^2}; ~~~~ cos(x) = \frac{1-u^2}{1+u^2}; ~~~~ sin(x) = \frac{2u}{1+u^2}\\ \int \frac{(1+u^2)^{2015}}{2u^2(1-u^2)^{2014}}du - 4028 \int \frac{(1+u^2)^{2013}}{(1-u^2)^{2014}}du \]

OpenStudy (freckles):

I think it is B. We can do small numbers and try to find a pattern. For example: I think \[\int\limits_{}^{}\frac{\csc^2(x)-n}{\cos^n(x)} dx=-\sec^{n}(x)\cot(x)+C ,\text{ where } n \text{ is a natural number }\] We can try to prove this by induction.

OpenStudy (freckles):

lets do where n is a natural number greater than 2

OpenStudy (freckles):

so starting at n=2 \[\int\limits_{}^{}\frac{\csc^2(x)-2}{\cos^2(x)} dx=\int\limits_{}^{}\frac{(\frac{u^2+1}{2u})^2-2}{(\frac{1-u^2}{1+u^2})^2} 2(u^2+1) du \\ \int\limits_{}^{}\frac{\frac{(u^2+1)^2}{4u^2}-2}{\frac{(1-u^2)^2}{(1+u^2)^2}} 2(u^2+1)du \\ \int\limits_{}^{} \frac{\frac{(u^2+1)^3}{2u^2}-4(u^2+1)}{\frac{(1-u^2)^2}{(1+u^2)^2} }du=\int\limits_{}^{}\frac{(u^2+1)^5-8u^2(u^2+1)^3}{2u^2 (1-u^2)^2} du\]

OpenStudy (freckles):

not finished

OpenStudy (freckles):

of course we could have done the other approach without this sub

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{(u^2)^5+5(u^2)^4+10(u^2)^3+10(u^2)^2+5(u^2)+1-8u^2((u^2)^3+3(u^2)^2+3(u^2)+1)}{2u^2(1-2u^2+u^4)}du\] \[\int\limits_{}^{}\frac{u^{10}+5u^8+10u^6+10u^4+5u^2+1-8u^8-24u^6-24u^4-8u^2}{2u^2(u^4-2u^2+1)} du\] \[\int\limits_{}^{}\frac{u^{10}-3u^8-14u^6-14u^4-3u^2+1}{2u^6-4u^4+2u^2} du\]

OpenStudy (freckles):

division...

OpenStudy (freckles):

Way too much typing...

OpenStudy (freckles):

this problem is crazy. where did it come from?

OpenStudy (freckles):

I will do it the other way...

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{\csc^2(x)-2}{\cos^2(x)}dx=\int\limits_{}^{}\frac{\csc^2(x)}{\cos^2(x)} dx -\int\limits_{}^{}\frac{2}{\cos^2(x)} dx \\ =\int\limits_{}^{}\frac{1}{\sin^2(x) \cos^2(x)} dx-\int\limits_{}^{}2 \sec^2(x) dx \\ =\int\limits_{}^{}\frac{\sin^2(x)+\cos^2(x)}{\sin^2(x)\cos^2(x)} dx-2 \tan(x)+C \] \[=\int\limits_{}^{}\sec^2(x) dx+\int\limits_{}^{}\csc^2(x) dx-2 \tan(x)+C\] \[=\tan(x)-\cot(x)-2\tan(x)+C=-\tan(x)-\cot(x)+C=\frac{-\sin(x)}{\cos(x)}+\frac{-\cos(x)}{\sin(x)}+C \\ = \frac{-\sin^2(x)-\cos^2(x)}{\cos(x)\sin(x)} +C =\frac{-1}{\cos(x)\sin(x)}+C=-\frac{\cos(x)}{\cos^2(x) \sin(x)}+C \\ =-\sec^2(x) \cot(x) +C\]

OpenStudy (freckles):

yeah that was much easier

OpenStudy (freckles):

for we just proved the first case

OpenStudy (freckles):

now if we assume \[\int\limits_{}^{}\frac{\csc^2(x)-k}{\cos^{k}(x)} dx=-\sec^{k}(x) \cot(x) +C \text{ for some integer } k \ge 2 \] can we show \[\int\limits_{}^{}\frac{\csc^2(x)-k-1}{\cos^{k+1}(x) } dx = -\sec^{k+1} \cot(x) +C \] let's play and try to show this. \[\int\limits_{}^{}\frac{\csc^2(x)-k}{\cos^{k+1}(x)} dx -\int\limits_{}^{} \frac{1}{\cos^{k+1}(x)} dx\] \[=\int\limits_{}^{}\frac{\csc^2(x)-k}{\cos^k(x)}\sec(x)dx-\int\limits_{}^{}\sec^{k+1} (x) dx \\ =-\sec^k(x)\cot(x)\sec(x)+\int\limits_{}^{}\sec^k(x)\cot(x)\sec(x)\tan(x) dx-\int\limits_{}^{}\sec^{k+1}(x) dx \\ =-\sec^{k+1}(x)\cot(x)+\int\limits_{}^{}\sec^{k+1} dx-\int\limits_{}^{}\sec^{k+1}(x) dx+C \\ =-\sec^{k+1}(x)\cot(x)+C \]

OpenStudy (freckles):

Therefore we have shown \[\int\limits_{}^{}\frac{\csc^2(x)-n}{\cos^n(x)} dx =-\sec^n(x) \cot(x)+C \text{ for all natural } n \ge 2 \]

OpenStudy (freckles):

by way i did use integration by parts in my inductive step

OpenStudy (ranga):

Thank you @freckles. Very nicely done! This was in one of the tests for a college entrance examination that my nephew in India took last year.

OpenStudy (freckles):

It is a tough question for an entrance exam I would think. But yeah I think the best thing to do here is start with small numbers and see what the pattern is.

OpenStudy (ranga):

Wolfram was choking not only on this integration but also on differentiating 3 out of the 4 answer choices! There is another interesting non-calculus question that was also asked in the same entrance exam which i will post next.

OpenStudy (freckles):

I will totally take a look.

OpenStudy (freckles):

omg ranga I just realized something we made that previous problem harder than it was I think...

OpenStudy (freckles):

\[\int\limits_{}^{} \frac{\csc^2(x)-2014}{\cos^{2014}(x)} dx \\ =\int\limits_{}^{}(\csc^2(x)-2014) \sec^{2014}(x) dx \\ \] \[=\int\limits_{}^{}\csc^2(x) \sec^{2014}(x) dx- \int\limits_{}^{}2014 \sec^{2014}(x) dx \\ = \text{ ( integration by parts ) } \\ -\cot(x)\sec^{2014}(x)+\int\limits_{}^{}\cot(x)2014\sec^{2013}(x)\sec(x)\tan(x) dx-\int\limits_{}^{}2014 \sec^{2014}(x) dx\] \[=-\cot(x)\sec^{2014}(x)+\int\limits_{}^{}2014\sec^{2014}(x)dx-\int\limits 2014\sec^{2014}(x) dx+C\]

OpenStudy (freckles):

I feel so dumb I didn't think of this earlier

OpenStudy (freckles):

giving us the answer we wanted which was \[=-\cot(x)\sec^{2014}(x)+C \]

OpenStudy (freckles):

it was just a simple integration by parts question

OpenStudy (ranga):

OMG! That is so cool! Thanks again.

OpenStudy (freckles):

@ranga I can't believe wolfram couldn't do that.

OpenStudy (freckles):

that is an awesome question lol

OpenStudy (ranga):

This test is full of such awesome questions. Not too much labor intensive but real good, thinking type of questions. I can go on and on. The ones my nephew got stuck in I do it under non-test condition at my own leisurely pace and solve a majority of them. But some I get stuck and I don't want to spend too much time on just one problem. BTW, I just posted the other algebra question from the same test.